# Constant Rate Scheduling: Constant-Rate Distributional Change for Efficient Training and Sampling in Diffusion Models

### v3

@@ Appendix E Implementation Details
E.1 Unet Model
The hyperparameters of the U-Net model are listed in Table 8. For latent-space diffusion models, we used the same values as the ImageNet 64 $\times$ 64 in ADM [52]. For pixel-space diffusion models, we used the same values as LSUN 256 $\times$ 256 in ADM [52].
@@ E.1 Unet Model
E.2 Hyperparameters for Training
The hyperparameters for training the diffusion models are summarized in Table 9. We saved ten checkpoints for each model. The FID score with 10K samples was evaluated for all checkpoints, and the results are reported with the checkpoint with the best FID score.
@@ E.2 Hyperparameters for Training
E.3 Noise Schedule for Training and Sampling
@@
When using a CRS-based training schedule, a common hyperparameter setting of $(\xi,\alpha_{\max},\alpha_{\min})=(1.0,1.0,0.0)$ is used (details are described in Appendix C.2). When using a CRS-based sampling schedule, the default setting is $(\xi,\alpha_{\max},\alpha_{\min})=(1.0,1.0,0.01)$. An exception was made when using CRS-$v_{x}+v_{\mathrm{FID}}$, where the hyperparameters are tuned according to the procedure described in Appendix D. For the experiments using pretrained EDMs, we used $(\alpha_{\max},\alpha_{\min})=(0.999998,0.0125)$, which is equivalent to the boundary condition of EDM: $(\sigma_{\min},\sigma_{\max})=(0.002,80)$. The boundary condition of EDM is transformed to $\alpha_{\max}$ and $\alpha_{\min}$ using Eq. (16).
@@
Sampling $\left\{\alpha_{\mathrm{linear}}\left(\frac{t}{T_{s}}\right)\Bigl{|}t=0,...,T_{s}\Bigr{.}\right\}$
@@
$\left\{\tilde{\alpha}_{\mathrm{shifted}}\left(\frac{t}{T_{s}}\right)\Bigl{|}t=0,...,T_{s}\Bigr{.}\right\},$
@@
[…]
@@
Appendix F Detailed Results
We present the detailed experimental results and confirm the superiority of CRS for optimizing both training and sampling schedules.
@@ F.1 Results on Pixel-Space Diffusion Models
LSUN Horse 256 $\times$ 256: The dependence on training and sampling schedules at $\mathrm{NFE}=250$ are listed in Table 11. To verify the generality with respect to samplers, we used three different samplers: SDE-DPM-Solver++(2M), PNDM, and DPM-Solver++(2M). As shown in the rows shaded in gray, multiple evaluation metrics, including FID score, are consistently improved when CRS-$v_{x}$ is used for optimizing the training schedule, regardless of the sampler.
Table 12 presents the dependence on the sampling schedules at $\mathrm{NFE}=250$ and $\mathrm{NFE}=50$. When using SDE-DPM-Solver++(2M), CRS-$v_{x}+v_{\mathrm{FID}}$ improved upon all four evaluation metrics at $\mathrm{NFE}=250$, achieving a new state-of-the-art FID score. When using PNDM and DPM-Solver++(2M), there was little difference between sampling schedules at $\mathrm{NFE}=250$. However, at $\mathrm{NFE}=50$, the performance became more sensitive to the sampling schedule, and CRS-$v_{x}+v_{\mathrm{FID}}$ achieved the best FID score.
LSUN Bedroom 256 $\times$ 256: Since only the FID score is reported in the main part of this study, we include sFID, precision, and recall in Table 13 for reference.
@@ F.2 Results on Latent-Space Diffusion Models
LSUN Church 256 $\times$ 256: The dependence on training and sampling schedules at $\mathrm{NFE}=30$ are listed in Table 14. To verify the generality with respect to samplers, we used five different samplers: Stochastic DDIM, SDE-DPM-Solver++(2M), DDIM, PNDM, and DPM-Solver++(2M). As shown in the rows shaded in gray, FID scores are consistently improved when CRS-$v_{x}$ is used for optimizing the training schedule across all samplers except for SDE-DPM-Solver++(2M).
@@
The results of evaluating the dependence on the sampling schedules using the stochastic and deterministic samplers are presented in Table 15 and 16, respectively. CRS-$v_{x}$ or CRS-$v_{\varepsilon}$ achieved the best FID score in all cases except when using SDE-DPM-Solver++(2M) at $\mathrm{NFE}=20$, and multiple evaluation metrics are simultaneously improved in most cases.
@@
LSUN Bedroom 256 $\times$ 256: The results of evaluating the dependence on the sampling schedules using the stochastic and deterministic samplers are presented in Table 17 and 18, respectively. CRS-$v_{x}$ is used for optimizing the training schedule. CRS achieved the best FID score in all cases except when using SDE-DPM-Solver++(2M) at $\mathrm{NFE}=30$ and $\mathrm{NFE}=20$, and multiple evaluation metrics are simultaneously improved in most cases.
@@
FID scores evaluated at small NFEs are listed in Table 19. Compared with the results reported in DPM-Solver-v3 [26], the better FID scores are achieved by using CRS for optimizing both the training and sampling schedules at $\mathrm{NFE}=10$ and $\mathrm{NFE}=20$. FID, sFID, precision, and recall are listed in Table 20. CRS based noise schedules consistently outperform the linear schedule.
@@
ImageNet 256 $\times$ 256: We demonstrated the superiority of CRS in class-conditional image generation. The dependence on training and sampling schedules at $\mathrm{NFE}=30$ are listed in Table 21. As shown in the rows shaded in gray, CRS-$v_{\mathrm{FID}}$ was the best for optimizing training schedule for ImageNet, while CRS-$v_{x}$, which consistently showed strong performance on other datasets, was the second-best. The reason CRS-$v_{\mathrm{FID}}$ performed better on ImageNet may be attributed to the fact that the feature model used to compute FID is trained on ImageNet classification.
@@
The results of evaluating the dependence on the sampling schedules using the stochastic and deterministic samplers are presented in Tables 22 and 23, respectively. We used CRS-$v_{x}$ for optimizing training schedule, which outperformed conventional noise schedules with significant margin on ImageNet and exhibited the best performance across a wide range of datasets. CRS achieved the best FID score in all cases except when using PNDM at $\mathrm{NFE}=20$ and DPM-Solver++(2M) at $\mathrm{NFE}=30$, and multiple evaluation metrics are simultaneously improved in most cases.
F.3 Results on Pretrained EDM
FFHQ 64 $\times$ 64: The evaluation results are listed in Table 24. When using DPM-Solver++(2M), CRS-$v_{x}+v_{\mathrm{FID}}$ outperforms both EDM and AYS. Using iPNDM further improves FID scores, and CRS-$v_{x}+v_{\mathrm{FID}}$ consistently performs better than EDM.
CIFAR10 32 $\times$ 32: The evaluation results on class-conditional image generation are listed in Table 25. CRS-$v_{x}+v_{\mathrm{FID}}$ consistently outperforms EDM, and significant performance gains are observed at small NFEs.
@@ Appendix G Generated Samples
Generated images are shown in the following figures:
  • LSUN Horse 256 $\times$ 256 in the pixel-space diffusion model: Figs. 14 and 15.
  • LSUN Bedroom 256 $\times$ 256 in the pixel-space diffusion model: Figs. 16 and 17.
  • LSUN Church 256 $\times$ 256 in the latent-space diffusion model: Fig. 18.
  • LSUN Bedroom 256 $\times$ 256 in the latent-space diffusion model: Fig. 19.
  • ImageNet 256 $\times$ 256 in the latent-space diffusion model: Fig. 20.
All models are trained using CRS-$v_{x}$ for optimizing training schedule.
Table 11: Dependence on training and sampling schedules for LSUN Horse 256 $\times$ 256 in pixel-space diffusion models at $\mathrm{NFE}=250$. To verify generality with respect to samplers, we used three different samplers. Rows shaded in gray represent dependence on training schedules, while unshaded rows represent dependence on sampling schedules. Bold values indicate best training or sampling schedules for each metric. Note that "cosine" refers to shifted cosine in this table.
@@
Table 12: Dependence on sampling schedules for LSUN Horse 256 $\times$ 256 in pixel-space diffusion models. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
@@
Table 13: Evaluation results on LSUN Bedroom 256 $\times$ 256 in pixel-space diffusion model. CRS-$v_{x}$ was used for optimizing training schedule.
@@
Table 14: Dependence on training and sampling schedules for LSUN Church 256 $\times$ 256 in latent-space diffusion models at $\mathrm{NFE}=30$. To verify generality with respect to samplers, we used five different samplers. Rows shaded in gray represent dependence on training schedules, while unshaded rows represent dependence on sampling schedules. Bold values indicate best training or sampling schedules for each metric.
@@
Table 15: Dependence on sampling schedules for LSUN Church 256 $\times$ 256 in latent-space diffusion models with stochastic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
@@
Table 16: Dependence on sampling schedules for LSUN Church 256 $\times$ 256 in latent-space diffusion models with deterministic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
@@
Table 17: Dependence on sampling schedules for LSUN Bedroom 256 $\times$ 256 in latent-space diffusion models with stochastic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
@@
Table 18: Dependence on sampling schedules for LSUN Bedroom 256 $\times$ 256 in latent-space diffusion models with deterministic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
@@
Table 19: FID scores evaluated on LSUN Bedroom 256 $\times$ 256 in latent-space diffusion model at small NFEs. Bold values indicate best FID scores at each NFE. Results for rows other than "Ours" are taken from previously published work [26].
@@
Table 20: Detailed results evaluated on LSUN Bedroom 256 $\times$ 256 in latent-space diffusion model at small NFEs. Bold values indicate best sampling schedules for each metric.
@@
Table 21: Dependence on training and sampling schedules for ImageNet 256 $\times$ 256 in latent-space diffusion models at $\mathrm{NFE}=30$. To verify generality with respect to samplers, we used five different samplers. Rows shaded in gray represent dependence on training schedule, while unshaded rows represent dependence on sampling schedules. Bold values indicate best training or sampling schedules for each metric.
@@
Table 22: Dependence on sampling schedules for ImageNet 256 $\times$ 256 in latent-space diffusion models with stochastic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.
@@
Table 23: Dependence on sampling schedules for ImageNet 256 $\times$ 256 in latent-space diffusion models with deterministic samplers. CRS-$v_{x}$ was used for optimizing training schedule. Bold values indicate best sampling schedules for each metric.

### v4

@@ Appendix E Implementation Details
E.1 U-Net Architecture
The hyperparameters of the U-Net model are listed in Table 13. For latent-space diffusion models, we use the same settings as those for ImageNet $\times64$ in ADM (Dhariwal and Nichol, 2021). For pixel-space diffusion models, we adopt the same settings as those for LSUN $\times256$ in ADM.
@@ E.1 U-Net Architecture
E.2 Training Hyperparameters
The training hyperparameters are summarized in Table 14. We save ten checkpoints for each model. FID scores are computed using 10K samples for all checkpoints, and we report the results from the checkpoint with the best FID.
@@ E.2 Training Hyperparameters
The training times for each model configuration are provided in Table 15. When using CRS to optimize the training schedule, training time increases by 20–30% compared to conventional noise schedules, due to the overhead of computing $v_{x}(\alpha)$ during training. In practice, however, we observe that the noise schedule stabilizes early; for example, on LSUN Horse $\times256$, $v_{x}(\alpha)$ shows minimal change beyond 10 epochs. This suggests that freezing the schedule after a small number of epochs or reducing the frequency of $\overline{D_{x}^{2}}$ updates can mitigate training overhead without degrading performance. Developing a more efficient implementation remains an important direction for future work.
Table 15: Training time of diffusion models.
Diffusion type Dataset GPUs Epochs Training schedule Training time
Latent space LSUN Church A100 $\times$ 8 1000
Linear
VDM++
CRS-$v_{\mathrm{FID}}$
85 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
105 hours
LSUN Bedroom A100 $\times$ 8 100
CRS-$v_{\mathrm{FID}}$
190 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
245 hours
ImageNet A100 $\times$ 8 500
Linear
VDM++
CRS-$v_{\mathrm{FID}}$
430 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
525 hours
Pixel space LSUN Horse H100 $\times$ 8 50
Linear
Shifted cosine
VDM++
170 hours
CRS-$v_{\varepsilon}$
CRS-$v_{x}$
215 hours
LSUN Bedroom H100 $\times$ 8 50
CRS-$v_{x}$
325 hours
E.3 Noise Schedules for Training and Sampling
@@
[…]
@@
[…]
@@
$\omega(\lambda)=\begin{cases}\underset{\lambda}{\operatorname{max}}\ \tilde{\omega}(\lambda),&\lambda<\underset{\lambda}{\operatorname{argmax}}\ \tilde{\omega}(\lambda)\\ \tilde{\omega}(\lambda),&\lambda\geq\underset{\lambda}{\operatorname{argmax}}\ \tilde{\omega}(\lambda),\end{cases},$
@@
Appendix F Detailed Experimental Results
We provide detailed experimental results to confirm the effectiveness of CRS for optimizing both training and sampling schedules.
@@ F.1 Results on Pixel-Space Diffusion Models
LSUN Horse $\times256$: The results showing the dependence on training and sampling schedules at $\mathrm{NFE}=250$ are presented in Table 17. To assess the generality across samplers, we evaluated three different ones: SDE-DPM-Solver++(2M), PNDM, and DPM-Solver++(2M). As shown in the gray-shaded rows, using CRS-$v_{x}$ for training consistently improved multiple evaluation metrics, including FID, regardless of the sampler used.
Figure 13 shows the evolution of the FID score with 10k samples during training, evaluated using DPM-Solver++ (2M) with $\mathrm{NFE}=250$. In this experiment, we compare CRS-$v_{x}$ with shifted cosine, which was the strongest-performing baseline training schedule. The results demonstrate that using CRS-$v_{x}$ as the training schedule not only improves the final FID score but also leads to faster convergence throughout training.
Table 18 presents the effect of different sampling schedules at $\mathrm{NFE}=250$ and $\mathrm{NFE}=50$. When using SDE-DPM-Solver++(2M), CRS-$v_{x}+v_{\mathrm{FID}}$ improved all four metrics at $\mathrm{NFE}=250$, achieving a new state-of-the-art FID score. With PNDM and DPM-Solver++(2M), differences among sampling schedules were minor at $\mathrm{NFE}=250$; however, at $\mathrm{NFE}=50$, performance became more sensitive to the sampling schedule. In this low-NFE regime, CRS-$v_{x}+v_{\mathrm{FID}}$ achieved the best FID score.
LSUN Bedroom $\times256$: Since only FID scores are reported in the main text (Table 2), we include sFID, precision, and recall in Table 19 for reference.
Table 20 compares the FID scores of CRS-$v_{x}+v_{\mathrm{FID}}$ and EDM sampling under various values of the hyperparameter $\rho$. While the default setting for EDM uses $\rho=7$, tuning $\rho$ can further improve its performance. However, even with the optimal $\rho$, EDM sampling did not outperform CRS-$v_{x}+v_{\mathrm{FID}}$ in terms of FID.
Figure 12: Evolution of the FID score with 10k samples during training on LSUN Horse $\times256$ in the pixel-space diffusion model. Using CRS-$v_{x}$ as the training schedule not only improves the final FID score but also leads to faster convergence compared to shifted cosine. Evaluation is performed using DPM-Solver++(2M) with $\mathrm{NFE}=250$.
Figure 13: Evolution of the FID score with 10k samples during training on LSUN Church $\times256$ in the latent-space diffusion model. CRS-$v_{x}$ achieves a slightly better best FID score and consistently lower FID scores before reaching the best value compared to the linear schedule. Evaluation is performed using DPM-Solver++(2M) with $\mathrm{NFE}=30$.
@@ F.2 Results on Latent-Space Diffusion Models
LSUN Church $\times256$: The results showing the dependence on training and sampling schedules at $\mathrm{NFE}=30$, using stochastic and deterministic samplers, are summarized in Tables 21 and 22, respectively. To evaluate generality across samplers, we tested five different ones: Stochastic DDIM, SDE-DPM-Solver++(2M), DDIM, PNDM, and DPM-Solver++(2M). As shown in the gray-shaded rows, CRS-$v_{x}$ consistently improved FID scores across all samplers except for SDE-DPM-Solver++(2M).
Figure 13 shows the evolution of the FID score with 10k samples during training, evaluated using DPM-Solver++ (2M) with $\mathrm{NFE}=30$. In this experiment, we compare CRS-$v_{x}$ with linear, which was the strongest-performing baseline training schedule. Both training schedules reach their best FID scores at around 400 epochs, with CRS-$v_{x}$ achieving a slightly better best FID score. Furthermore, CRS-$v_{x}$ also showed consistently lower FID scores before reaching the best value, indicating its effectiveness in terms of convergence speed as well.
@@
The effects of different sampling schedules are shown separately for stochastic and deterministic samplers in Tables 23 and 24, respectively. CRS-$v_{x}$ or CRS-$v_{\varepsilon}$ achieved the best FID scores in all cases except when using SDE-DPM-Solver++(2M) at $\mathrm{NFE}=20$. In most cases, multiple evaluation metrics—including sFID, precision, and recall—were simultaneously improved.
@@
LSUN Bedroom $\times256$: The results showing the dependence on sampling schedules using stochastic and deterministic samplers are presented in Tables 25 and 26, respectively. The training schedule was fixed to $v_{x}. CRS$ achieved the best FID scores in all cases, except when using SDE-DPM-Solver++(2M) at $\mathrm{NFE}=30$ and $\mathrm{NFE}=20$. In most cases, multiple evaluation metrics were simultaneously improved.
@@
FID scores evaluated at small NFE settings are listed in Table 27. Compared to the results reported in DPM-Solver-v3 (Zheng et al., 2023b), better FID scores are consistently achieved by using CRS to optimize both training and sampling schedules across all NFE values. As with other evaluations in this paper, we set $\xi=1$ by default for CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, except at $\mathrm{NFE}=5$, where $\xi$ was manually tuned to further improve performance. In this experiment, we observed that the importance of tuning $\xi$ increases as the NFE decreases. We attribute this behavior to the discretization error introduced by the continuous-time approximation used in Eq. (7). A comprehensive set of evaluation metrics, including FID, sFID, precision, and recall, is reported in Table 28. CRS-based noise schedules consistently outperform the linear schedule across all metrics, except for recall at $\mathrm{NFE}=20$.
ImageNet $\times256$: We demonstrated the effectiveness of CRS in class-conditional image generation on ImageNet. It is important to note that our experiments do not use classifier-free guidance (CFG), and therefore the FID scores are naturally higher than those obtained under CFG-enabled settings. However, as shown in Table 29, our model achieves better FID score than the LDM-4 results reported in the original LDM paper (Rombach et al., 2022). These observations indicate that our ImageNet results represent a reasonable level of performance for class-conditional generation without CFG, and provide a meaningful basis for evaluating the impact of CRS in this setting.
@@
The results showing the dependence on training and sampling schedules at $\mathrm{NFE}=30$, using both stochastic and deterministic samplers, are summarized in Tables 30 and 31, respectively. As shown in the gray-shaded rows, CRS-$v_{\mathrm{FID}}$ achieved the best performance for training schedule optimization, while CRS-$v_{x}$, which consistently performed well on other datasets, was the second-best. This result may be attributed to the fact that the feature extractor used in FID computation is pre-trained on ImageNet, which may advantage CRS-$v_{\mathrm{FID}}$ in this specific setting.
@@
The impact of different sampling schedules using stochastic and deterministic samplers is shown in Tables 32 and 33, respectively. Here, the training schedule was fixed to CRS-$v_{x}$, which outperformed conventional noise schedules by a significant margin on ImageNet and delivered strong results across a wide range of datasets. CRS achieved the best FID scores in all cases, except when using PNDM at $\mathrm{NFE}=20$ and DPM-Solver++(2M) at $\mathrm{NFE}=30$. In most settings, CRS-based schedules also led to improvements in other evaluation metrics, including sFID, precision, and recall.
@@ Appendix G Generated Samples
Generated samples are shown in the following figures:
  • LSUN Horse $\times256$ in the pixel-space diffusion model: Figs. 14 and 15.
  • LSUN Bedroom $\times256$ in the pixel-space diffusion model: Figs. 16 and 17.
  • LSUN Church $\times256$ in the latent-space diffusion model: Fig. 18.
  • LSUN Bedroom $\times256$ in the latent-space diffusion model: Fig. 19.
  • ImageNet $\times256$ in the latent-space diffusion model: Fig. 20.
All models are trained using CRS-$v_{x}$ for the training schedule optimization.
Table 17: Performance of pixel-space diffusion models on LSUN Horse $\times256$ at $\mathrm{NFE}=250$, evaluated under various training and sampling schedules. To assess generality across samplers, we used three different samplers: SDE-DPM-Solver++(2M), PNDM, and DPM-Solver++(2M). Gray-shaded rows indicate results for different training schedules (with fixed sampling), while unshaded rows show results for different sampling schedules (with fixed training). Bold values indicate the best performance for each metric. Note that "cosine" refers to the shifted cosine schedule.
@@
Table 18: Performance of pixel-space diffusion models on LSUN Horse $\times256$ under different sampling schedules. Training schedule was fixed to CRS-$v_{x}$. Bold values indicate the best sampling schedule for each evaluation metric.
@@
Table 19: Evaluation results on LSUN Bedroom $\times256$ in pixel-space diffusion model. CRS-$v_{x}$ was used for optimizing training schedule.
@@
Table 20: Comparison of FID scores for CRS-$v_{x}+v_{\mathrm{FID}}$ and EDM sampling on LSUN Bedroom $\times256$ in the pixel-space diffusion model. Performance of EDM is reported under various values of the hyperparameter $\rho$. Even with the optimal $\rho$, EDM does not outperform CRS-$v_{x}+v_{\mathrm{FID}}$.
Sampler Sampling schedule $\mathrm{NFE}=5$ $\mathrm{NFE}=10$
DPM-Solver++(2M) EDM: $\rho=1$ 190.32 102.10
EDM: $\rho=3$ 30.24 13.12
EDM: $\rho=5$ 22.97 7.92
EDM: $\rho=7$ 23.72 4.73
CRS-$v_{x}+v_{\mathrm{FID}}$ 14.02 4.88
UniPC EDM: $\rho=1$ 178.24 90.30
EDM: $\rho=3$ 23.45 4.43
EDM: $\rho=5$ 55.06 3.96
EDM: $\rho=7$ 82.52 4.94
CRS-$v_{x}+v_{\mathrm{FID}}$ 21.42 3.30
Table 21: Performance of latent-space diffusion models on LSUN Church $\times256$ at $\mathrm{NFE}=30$, evaluated under various training and sampling schedules using stochastic samplers. Two samplers were used: Stochastic DDIM and SDE-DPM-Solver++(2M). Gray-shaded rows indicate variations in training schedules (with fixed sampling), while unshaded rows show variations in sampling schedules (with fixed training). Bold values indicate the best result for each evaluation metric.
@@
Table 22: Performance of latent-space diffusion models on LSUN Church $\times256$ at $\mathrm{NFE}=30$, evaluated under various training and sampling schedules using deterministic samplers. Three samplers were used: DDIM, PNDM, and DPM-Solver++(2M). Gray-shaded rows indicate variations in training schedules (with fixed sampling), while unshaded rows show variations in sampling schedules (with fixed training). Bold values indicate the best result for each evaluation metric.
Training Sampling Metrics
@@
Table 23: Performance of latent-space diffusion models on LSUN Church $\times256$ using stochastic samplers, evaluated under different sampling schedules. The training schedule was fixed to CRS-$v_{x}$. Bold values indicate the best sampling schedule for each evaluation metric.
@@
Table 24: Performance of latent-space diffusion models on LSUN Church $\times256$ using deterministic samplers, evaluated under different sampling schedules. The training schedule was fixed to CRS-$v_{x}$. Bold values indicate the best sampling schedule for each evaluation metric.
@@
Table 25: Performance of latent-space diffusion models on LSUN Bedroom $\times256$ using stochastic samplers, evaluated under different sampling schedules. The training schedule was fixed to CRS-$v_{x}$. Bold values indicate the best sampling schedule for each evaluation metric.
@@
Table 26: Performance of latent-space diffusion models on LSUN Bedroom $\times256$ using deterministic samplers, evaluated under different sampling schedules. The training schedule was fixed to CRS-$v_{x}$. Bold values indicate the best sampling schedule for each evaluation metric.
@@
Table 27: FID scores on LSUN Bedroom $\times256$ in latent-space diffusion models at small NFE settings. Bold values indicate the best FID score at each NFE. Results for rows other than "Ours" are taken from Zheng et al. (2023b). As with other evaluations in this paper, we set $\xi=1$ by default for CRS-$v_{x}$ and CRS-$v_{\varepsilon}$, except at $\mathrm{NFE}=5$, where $\xi$ was manually tuned to further improve performance. In this experiment, we observed that the importance of tuning $\xi$ increases as the NFE decreases. We attribute this behavior to the discretization error introduced by the continuous-time approximation used in Eq. (7).
@@
CRS-$v_{\varepsilon}$: $\xi=1.4$ 7.64
CRS-$v_{x}$: $\xi=2.2$ 6.96
Table 28: Detailed evaluation results on LSUN Bedroom $\times256$ in latent-space diffusion models at small NFE settings. Bold values indicate the best sampling schedule for each metric.
@@
CRS-$v_{\varepsilon}$: $\xi=1.4$ 7.64 9.23 0.45 0.52
CRS-$v_{x}$: $\xi=2.2$ 6.96 8.65 0.45 0.49
Table 29: Comparison between our trained model and LDM-4 (Rombach et al., 2022) on ImageNet under class-conditional generation without CFG. Nparams denotes the number of parameters in the U-Net model. Our model, trained with CRS-$v_{x}$ as the training schedule, achieves a better FID score than LDM-4 despite having fewer U-Net parameters.
Model Training schedule Sampling schedule Sampler NFE FID($\downarrow$) Nparams
LDM-4 Linear Linear Stochastic DDIM 250 10.56 400M
Ours CRS-$v_{x}$ 9.67 296M
Table 30: Performance of latent-space diffusion models on ImageNet $\times256$ at $\mathrm{NFE}=30$, evaluated under various training and sampling schedules using stochastic samplers. Two samplers were used: Stochastic DDIM and SDE-DPM-Solver++(2M). Gray-shaded rows indicate variations in training schedules (with fixed sampling), while unshaded rows show variations in sampling schedules (with fixed training). Bold values indicate the best result for each evaluation metric.
@@
Table 31: Performance of latent-space diffusion models on ImageNet $\times256$ at $\mathrm{NFE}=30$, evaluated under various training and sampling schedules using deterministic samplers. Three samplers were used: DDIM, PNDM, and DPM-Solver++(2M). Gray-shaded rows indicate variations in training schedules (with fixed sampling), while unshaded rows show variations in sampling schedules (with fixed training). Bold values indicate the best result for each evaluation metric.
Training Sampling Metrics
@@
Table 32: Performance of latent-space diffusion models on ImageNet $\times256$ using stochastic samplers, evaluated under different sampling schedules. The training schedule was fixed to CRS-$v_{x}$. Bold values indicate the best sampling schedule for each evaluation metric.
@@
Table 33: Performance of latent-space diffusion models on ImageNet $\times256$ using deterministic samplers, evaluated under different sampling schedules. The training schedule was fixed to CRS-$v_{x}$. Bold values indicate the best sampling schedule for each evaluation metric.